OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (-4, 0) and a directrix at x = 4.

OpenStudy (anonymous):

@mathmale

hero (hero):

@griffin5210, are you still here?

OpenStudy (anonymous):

yes i need help @Hero

hero (hero):

In general, if you are given two points, the focus \((x_1, y_1)\) and the directrix \((x_2, y_2)\), then you can insert the coordinates of both points in to the distance formula: \((x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\) In this case: \((x_1, y_1) = (-4,0)\) \((x_2, y_2) = (4, y)\)